johnnyappleseed:

This year, the selling price of each calendar was $13.25. The price this year represents 6% more than the selling price of each calendar last year. What was the selling price of each calendar last year?

JamesTDG:

Steps to solve: 6% = 0.06, 0.06*13.25 = (6*13.225)/100 = 79.5/100 = 0.795, round to nearest 100th, 0.80. 13.25 - 0.80 = 12.45

Florisalreadytaken:

how did u come up with that? anyway, it says that: `The price this year represents 6% more than the selling price of each calendar last year.` you knwow that \( 6\% = \frac{5}{100} \) right? we can find the price of the calendars from net year by taking \( \frac{6}{100} \) off the \( \frac{100}{100} \) so, \[ \frac{100}{100} - \frac{6}{100} = \frac{94}{100} \] we know that \( \frac{94}{100} = 0.94 \) use the calculator, and do this \[ 13.25 \ \cdot \ 0.94 = \text{our answer} \]

JamesTDG:

You made a little typo there...

Florisalreadytaken:

yeah i noticed -- i was rushing!